# Flexible MOF Generation with Torsion-Aware Flow Matching

### v3

@@ G.3 Energy-level evaluation using MLIP
We further evaluate the energy levels of generated MOFs using UMA [52], a state-of-the-art machine-learned interatomic potential. Specifically, we generate 10,000 structures with both MOFFlow-2 and MOFDiff and compute the energy per atom (eV/atom) using the uma-m-1.1 model. As shown in Figure˜9(b), the energy distribution of structures generated by MOFFlow-2 aligns more closely with the training dataset than that of MOFDiff.
@@ Appendix H Limitations
Although MOFFlow-2 demonstrates strong potential for MOF structure prediction and generation, several limitations remain. Firstly, because the pipeline relies on RDKit for initial conformer generation, it is challenging to predict structures whose organic building blocks are chemically invalid or incomplete; for example, those lacking carboxylate groups, which are common in MOF decomposition schemes [37]. Second, for conditional generation in Section˜G.2, only the building block generator is conditioned on the property, whereas the structure prediction module may also depend on it. Finally, our evaluation of energy levels with machine learning interatomic potentials is less accurate than that with density functional theory (DFT).

### v4

@@ G.3 Energy-level evaluation using MLIP
We further evaluate the energy levels of generated MOFs using UMA [wood2025family], a state-of-the-art machine-learned interatomic potential. Specifically, we generate 10,000 structures with both MOFFlow-2 and MOFDiff and compute the energy per atom (eV/atom) using the uma-m-1.1 model. As shown in Figure˜9(b), the energy distribution of structures generated by MOFFlow-2 aligns more closely with the training dataset than that of MOFDiff.
@@ Appendix H Limitations
Although MOFFlow-2 demonstrates strong potential for MOF structure prediction and generation, several limitations remain. Firstly, because the pipeline relies on RDKit for initial conformer generation, it is challenging to predict structures whose organic building blocks are chemically invalid or incomplete; for example, those lacking carboxylate groups, which are common in MOF decomposition schemes [bucior2019identification]. Second, for conditional generation in Section˜G.2, only the building block generator is conditioned on the property, whereas the structure prediction module may also depend on it. Finally, our evaluation of energy levels with machine learning interatomic potentials is less accurate than that with density functional theory (DFT).